OpenStudy (anonymous):

does f(x) = x^3 - 1 + sin(x) have a root in (0,1)? Explain

OpenStudy (anonymous):

no??

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

thank-you so much! im suppose to justify each step. doesn't feel like there's much to justify.

OpenStudy (anonymous):

basically says: Explain your answer carefully, and be sure to justify each assertion you make.

OpenStudy (anonymous):

seems so unnecessary

OpenStudy (anonymous):

haha, I agree. Do you have time to help me with 1 more q?

OpenStudy (anonymous):

let a and be denote constants. find the values of a and b so that the function (1+e^-x if x<0 and ax+b if x>or=0) is differentiable everywhere

OpenStudy (anonymous):

"and b"

OpenStudy (anonymous):

im following

OpenStudy (anonymous):

when i solve for a, I will also have x

OpenStudy (anonymous):

how will i isolate a

OpenStudy (anonymous):

x could be equal to greater than 0?! so I could set x as 1, when solving for a?

OpenStudy (anonymous):

i am so wrong.

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

b=2; a=0

hartnn (hartnn):

since the function is differentiable everywhere, f' (x) at x=0 exist so, f'(x) at x->0- = f'(x) at x->0+ d/dx (1+e^-x) = d/dx (ax+b) at x=0

hartnn (hartnn):

you can get a from here

hartnn (hartnn):

tell me if you did not get what i did

OpenStudy (anonymous):

gotchya

hartnn (hartnn):

so what you get for a ?

OpenStudy (anonymous):

o :S

hartnn (hartnn):

derivative of 1+e^-x is -e^-x right ? when x=o, left side = -e^0 = -1 got this ?

OpenStudy (anonymous):

yes, got that

hartnn (hartnn):

right side ?

hartnn (hartnn):

derivative of ax+b is just a so, overall, a = -1 :)

OpenStudy (anonymous):

eeeeek, how come it's -?

OpenStudy (anonymous):

b=2; a=-1 then?

OpenStudy (anonymous):

comeee back, hartnnnn.

hartnn (hartnn):

yes, thats correct.

hartnn (hartnn):

you got the steps, right ? and how it is - ?

OpenStudy (phi):

hartnn answered your second question the idea is that the function 1+e^(-x) --> 2 as x-->0, i.e. the point (0,2) you want the second function, defined for x≥0 to start at (0,2) so that you do not have a discontinuity. in other words, you want a x + b= 2 when x is 0. that gives you b=2 you also want the function to be differentiable. That means you want the same differential for 1+e^(-x) as x-->0 as you get for a x + b at x=0 the differential of 1+e^(-x) = -e^(-x) (take the derivative) as x->0, -e^(-x) --> -1 one way to see this: \[ -e^{-x}= \frac{-1}{e^x} \\ \lim_{x->0} \frac{-1}{e^x}= \frac{-1}{1}= -1 \] that means you want the derivative of a x + b to match -1 when x is 0 d (ax+b)= a you want a= -1 the final answer is -x+2 here is a graph of the situation

OpenStudy (phi):

For your first question ***does f(x) = x^3 - 1 + sin(x) have a root in (0,1)? Explain*** the interval does not include 0 or 1 (the parens mean exclude the boundary values). However, if we evaluate f(x) at the boundaries we find f(0)= 0 - 1 + sin(0)= -1 f(1)= 1-1 +sin(1) = 0.84 (x is assumed to be in radians) by the intermediate value theorem http://www.mathsisfun.com/algebra/intermediate-value-theorem.html there must be some x within the closed interval [0+, 1-] where f(x)=0 (by [0+,1-] I mean the interval +episilon to 1 - epsilon, because I think the intermediate value theorem wants a closed interval.